OpenStudy (anonymous):

Integrate 3^sin(x) cosx

OpenStudy (tkhunny):

Have you considered u = sin(x)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i guess I got stuck at \[\int\limits_{?}^{?}3^u\]

OpenStudy (anonymous):

du

OpenStudy (tkhunny):

\(\int 3^{u}\;du = \dfrac{3^{u}}{ln(3)} + C\) That should be in a book, somewhere.

OpenStudy (anonymous):

replace 3^sin(x) with e^(ln(3)*sinx)

OpenStudy (anonymous):

and then use sub:u=ln(3)*sinx

OpenStudy (anonymous):

that seems more complicated than it has to be :) but ill try it

OpenStudy (tkhunny):

@ASAAD123 is demonstrating a good general method for all sorts of exponential things. It is a little more complicated than it needs to be, but only visually. If you can get past the eye pain, it should work. Or, you can just know about exponentials with bases other than e, which is where I went. Why do you care about the limits? Is this definite integral and you just didn't tell us? If so, you'll have to play with the limits, too.

OpenStudy (anonymous):

does the e^ln cancel?

OpenStudy (anonymous):

no it is easier ,\[\int\limits_{}^{}e ^{(\ln3) \times sinx}cosx dx\]

OpenStudy (anonymous):

u=ln3 *sinx du=ln3 *cosx dx ................

OpenStudy (anonymous):

so no product rule to find du?

OpenStudy (anonymous):

no ,why ? ln3 is a number.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

how do yall master this stuff? there's so many rules.

OpenStudy (tkhunny):

Please don't ask about cancelling. Exponents and logarithms are functions requiring arguments. \(e^{ln}\) doesn't mean ANYTHING. Like I said, it's fine if you can get past the visual pain. It is more visually complicated. There was no product rule int he first place. Find methods that work well for you and try to use the ones with which you are most comfortable. Also, and perhaps more importantly, keep everything in yer head and do some exploration. Do things the easy way and the hard way and any other way you can think of. You will gain experience.

OpenStudy (anonymous):

so my final answer is \[\frac{ 1 }{ \ln3 }e^{\ln3sinx}+c\] +C

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

or \[\frac{ 3^{sinx} }{ \ln 3 }+c\]

OpenStudy (anonymous):

I wont even ask how you got that or how it is the same as what I got lol

OpenStudy (anonymous):

oh it canceled, nvm I see lol

OpenStudy (anonymous):

TYsoVM